OpenStudy (anonymous):

I'm trying to find the derivative of -6/x at x = 12

OpenStudy (bibby):

\(\dfrac{d}{dx}cf(x)=c*f'(x)\) separate it as \(-6*\dfrac{1}{x}\)

OpenStudy (anonymous):

\[\frac{ \frac{ -6 }{ x-h } - \frac{ -6 }{ x } }{ h }\]

OpenStudy (anonymous):

I've never seen what you said before @bibby

OpenStudy (bgrg007):

you just derive 1/x, then multiply whatever you get by -6. substitute x for 12 and solve.

OpenStudy (cggurumanjunath):